# Double-coset zeta functions for groups acting on trees

### v1

@@ Abstract.
We study the double-coset zeta functions for groups acting on trees, focusing mainly on weakly locally $\infty$-transitive or (P)-closed actions. After giving a geometric characterisation of convergence for the defining series, we provide explicit determinant formulae for the relevant zeta functions in terms of local data of the action. Moreover, we prove that evaluation at $-1$ satisfies the expected identity with the Euler–Poincaré characteristic of the group. The behaviour at $-1$ also sheds light on a connection with the Ihara zeta function of a weighted graph introduced by A. Deitmar.
@@ 1. Introduction
Double cosets play a prominent role in multiple aspects of group theory and beyond. For instance, they are the building blocks of Hecke algebras. Regarded as collections of cosets, they describe spheres with respect to the Weyl distance in the building associated with a Bruhat decomposition of a group [1]. Strictly related objects, namely the suborbits, are also widely studied in permutation group theory. It is common to arrange the suborbit sizes – provided they are all finite – in a non-decreasing sequence $(a_{n})_{n\geq 1}$ and estimate, for example, the growth of $a_{n}$ as a function of $n$ [4] [21, §5].
@@
If for every $n\geq 1$, $a_{n}(G,H,K)$ is finite and polynomially bounded as a function in $n$, we say that $(G,H,K)$ has polynomial double-coset growth.
If $(G,H,K)$ has polynomial double-coset growth, we can consider the Dirichlet series generated by $(a_{n}(G,H,K))_{n\geq 1}$, that is,
@@
where $s$ is a complex variable. By [14, §I.1, Theorem 3], this series converges in some half-plane $\{s\in\mathbb{C}\mid\mathrm{Re}(s)>\alpha\}$ and the function it determines is called the double-coset zeta function of $(G,H,K)$.
Zeta functions initially arose in number theory, although nowadays they are also an established tool in studying groups, rings and algebras (see M. du Sautoy’s survey [19] for motivation). In group theory, the seminal work of F. Grunewald, D. Segal, G. Smith [13] initiated the study of numerous zeta functions associated to finitely generated nilpotent groups or profinite groups. Double-coset zeta functions have been one of the first instances of zeta functions in the class of totally disconnected locally compact (= t.d.l.c.) groups which are possibly neither discrete nor profinite. Their introduction was motivated by an interesting behaviour at $s=-1$. I. Castellano, G. Chinello and T. Weigel [8] have provided examples of unimodular t.d.l.c. groups $G$ with a compact open subgroup $K\leq G$ for which the meromorphic continuation of $\zeta_{G,K,K}(s)$, evaluated at $s=-1$, recovers the Euler–Poincaré characteristic $\widetilde{\chi}_{G}$ of $G$ (in the sense of [8, §5]). Namely, one has
@@
The connection between growth series and Euler–Poincaré characteristics is not an isolated phenomenon, see for instance [11, 12, 23] or the introduction of [8]. One of the main goals of the present paper is to prove that the Euler–Poincaré identity holds in two relevant classes of unimodular t.d.l.c. groups acting on trees. More generally, we present a systematic study of the double-coset zeta functions for groups (not necessarily t.d.l.c.) acting on trees, including convergence criteria and explicit formulae.
In what follows, every graph $\Gamma=V\Gamma\sqcup E\Gamma$ is meant in the sense of J-P. Serre [20], and $V\Gamma$ and $E\Gamma$ denote the set of vertices and the set of edges of $\Gamma$, respectively (cf. Section 2.2). Every group action on a tree is without inversion of edges and without global fixed points. The tree of the action is always leafless and with at least one edge (cf. Section 3.1). Moreover, at least for the theorems of the introduction, the stabilisers of adjacent vertices are assumed to be incomparable with respect to the inclusion.
@@
The first property is weak local $\infty$-transitivity. A group action on a tree $(G,T)$ is weakly locally $\infty$-transitive if, for every $v\in VT$ and every path $\mathfrak{p}$ in the quotient graph $G\backslash T$, the stabiliser of $v$ acts transitively on the set of all geodesics in $T$ starting at $v$ and lifting $\mathfrak{p}$. This condition, which we introduce in Section 3.5, generalises the well-known notion of local $\infty$-transitivity to non-edge transitive group actions.
The second property we consider is (P)-closedness. The concept of (P)-closed actions on trees has been introduced by C. Banks, M.  Elder and G. Willis [2, §3]. It stems from the slightly more general concept of actions on trees with Tits’ independence property, introduced by J. Tits [25, §4.2]. The latter properties play a central role in the theory of groups acting on trees and even beyond. For instance, one may use a (P)-closed action to produce simple groups acting on trees (cf. [25, Théorème 4.5], [2, Theorem 7.3] or [17, Theorem 1.8]). Remarkably, C. Reid and S. Smith [17] provide a complete classification of (P)-closed actions on trees by their local action diagram, which is a local datum attached to each group action on a tree. In Section 3.4 we briefly recall it. With this local description, one may study several global properties of the group action through more accessible features of the associated local action diagram. In the present paper, we exploit this approach more than once.
The reader may find it convenient to keep the following example in mind.
@@
Let $\Gamma$ be a connected graph with a function $\omega\colon E\Gamma\to\mathbb{Z}_{\geq 1}$ (called edge weight). Following [6, §3.5], the pair $(\Gamma,\omega)$ admits an essentially unique universal cover $(T,\pi)$, which consists of a tree $T$ and a graph epimorphism $\pi\colon T\to\Gamma$ with the following property: for all $a\in E\Gamma$ and $v\in VT$ with $\pi(v)=o(a)$, the number of edges $e\in ET$ with origin $v$ and $\pi(e)=a$ is exactly $\omega(a)$. Note that $\pi$ is not a covering map in the usual sense. Indeed, the degree $\mathrm{deg}_{T}(v)=|o^{-1}(v)|$ of every $v\in VT$ in $T$ might be greater than the degree $\mathrm{deg}_{\Gamma}(\pi(v))=|o^{-1}(\pi(v))|$ of $\pi(v)$ in $\Gamma$, namely
$\mathrm{deg}(v)=\sum_{a\in E\Gamma\,:\,o(a)=\pi(v)}\omega(a)\geq\mathrm{deg}_{\Gamma}(\pi(v)).$
The group of deck transformations of $(\Gamma,\omega)$ is
@@
where $\mathrm{Aut}(T)$ denotes the group of automorphisms of $T$. The group $\mathrm{Aut}_{\pi}(T)$ is t.d.l.c. with respect to the subspace topology induced by $\mathrm{Aut}(T)$. Moreover, the $\mathrm{Aut}_{\pi}(T)$-action on $T$ is both weakly locally $\infty$-transitive and (P)-closed (cf. Example 3.12(i) and [25, §4.2]).
@@
The first step towards the study of double-coset zeta functions is to determine under which conditions a triple $(G,H,K)$ has the double-coset property or polynomial double-coset growth. Here we only consider the case of $G$ acting on a locally finite tree and we choose $H,K$ among the stabilisers of vertices or edges of $T$ (written $H=G_{t_{1}}$ and $K=G_{t_{2}}$, for $t_{1},t_{2}\in T$). In this setting, we provide the following characterisation.
Theorem A (cf. Theorems 5.6 and 5.7).
Let $G$ be a group that acts on a locally finite tree $T$ with a finite quotient graph. Assume that the action $(G,T)$ is weakly locally $\infty$-transitive or (P)-closed. Then the following are equivalent, for all $t_{1},t_{2}\in T$:
@@
    there is $k\geq 1$ such that, for every geodesic $\mathfrak{p}$ in $T$ of length $l\geq 1$, the pointwise stabiliser of $\mathfrak{p}$ does not fix any geodesic in $T$ of length $l+k$ extending $\mathfrak{p}$.
Theorem A implies that both the double-coset property and the polynomial double-coset growth are independent from the choice of $t_{1}$ and $t_{2}$. A similar independence has been shown in [8, Proposition 6.2].
It is worth mentioning that the chain of implications (iii)$\Rightarrow$(ii)$\Rightarrow$(i) is true even when dropping the hypothesis of weak local $\infty$-transitivity or (P)-closedness (cf. Proposition 5.5).
@@
In the weakly locally $\infty$-transitive case, counting the $(H,K)$-double cosets is rephrased in a more accessible counting of certain paths in the quotient graph $\Gamma$ with a suitably defined weight (cf. Section 3.2). This also suggests the definition of a more general Dirichlet series $\mathcal{Z}_{\Gamma,u_{1}\to u_{2}}(s)$ associated to an arbitrary graph $\Gamma$ with edge weight $\omega\colon E\Gamma\to\mathbb{Z}_{\geq 1}$ and to $u_{1},u_{2}\in\Gamma$ (cf. Definition 5.9). The series $\mathcal{Z}_{\Gamma,u_{1}\to u_{2}}(s)$ recovers $\zeta_{G,G_{t_{1}},G_{t_{2}}}(s)$ whenever $G$ is a group acting weakly locally $\infty$-transitively on a locally finite tree $T$ with quotient graph $\Gamma$ and standard edge weight $\omega$ (cf. Section 3.2), and whenever $t_{1},t_{2}\in T$ satisfy $G\cdot t_{1}=u_{1}$ and $G\cdot t_{2}=u_{2}$.
@@
In both cases, we can borrow ideas and techniques from graph theory (e.g., counting paths in graphs by their weight, see for instance [9] or [10]) and provide explicit formulae for $\zeta_{G,G_{t_{1}},G_{t_{2}}}(s)$, for all $t_{1},t_{2}\in T$.
@@
Theorem B (cf. Theorems 5.12 and 5.19).
Let $(G,T)$ be a group action on a tree that is weakly locally $\infty$-transitive or (P)-closed. Let $t_{1},t_{2}\in T$ be such that $(G,G_{t_{1}},G_{t_{2}})$ has polynomial double-coset growth. Then,
@@
for explicitly defined square matrices $\mathcal{E}^{\bullet}(s)$ and $\mathcal{U}^{\bullet}_{t_{1},t_{2}}(s)$ whose entries are entire functions in $s\in\mathbb{C}$, and for a determined integer $\epsilon^{\bullet}_{t_{1}}(t_{2})$. Here $I^{\bullet}$ denotes the identity matrix of the same dimension of $\mathcal{E}^{\bullet}(s)$. In particular, $\zeta^{\bullet}_{G,G_{t_{1}},G_{t_{2}}}(s)$ extends to a meromorphic function over $\mathbb{C}$.
In Theorem B, the matrix $\mathcal{E}^{\bullet}(s)$ can be interpreted as a weighted adjacency matrix of the local structure in which we count the paths (cf. Definitions 5.10 and 5.17). The matrix $\mathcal{U}_{t_{1},t_{2}}^{\bullet}(s)$ and the integer $\epsilon^{\bullet}_{t_{1}}(t_{2})$ can be regarded as “perturbation data” given by the choice of $t_{1}$ and $t_{2}$.
@@
The behaviour at $s=-1$ and the Euler–Poincaré characteristic
One of the main goals of the paper is the study of the local behaviour at $s=-1$ of the relevant double-coset zeta functions. Unless it is a finite sum, the Dirichlet series in (1.2) does not converge at $s=-1$. This underlines the importance of having a continuation of $\zeta_{G,H,K}(s)$ at least to $s=-1$ – which in our context is provided by Theorem B – to carry out such an evaluation.
Addressing [8, Question G(b)], we prove that the Euler–Poincaré identity also holds in our framework. A crucial step towards this goal is to reduce the evaluation at $s=-1$ to the more accessible weakly locally $\infty$-transitive case (cf. Lemma 7.10). More specifically, we can only focus on $\mathcal{Z}_{\Gamma,u\to u}(-1)^{-1}$ for a finite connected edge-weighted graph $(\Gamma,\omega)$. After introducing suitable notions of unimodularity and Euler–Poincaré characteristic $\chi(\Gamma,u)$ at $u\in\Gamma$ on $(\Gamma,\omega)$ (cf. Definition 7.4), we deduce the following.
@@
In Theorem E, the hypothesis on $\omega$ guarantees that a formula analogous to the one in Theorem B is applicable to $\mathcal{Z}_{\Gamma,u\to u}(s)$, for all $u\in\Gamma$. The proof of Theorem E is based on some splitting formulae satisfied by $\mathcal{Z}_{\Gamma,u\to u}(s)^{-1}$, $u\in\Gamma$, that we discuss in Section 6. The assumption that $\Gamma$ has no cycles of length $\geq 2$ guarantees their general applicability. By Theorem E and Lemma 7.10, we deduce the following.
@@
Let $G$ be a unimodular t.d.l.c. group acting on a locally finite tree $T$ with compact open vertex stabilisers. Assume that the quotient graph is finite and does not have cycles of length $\geq 2$. Suppose also that $(G,T)$ is weakly locally $\infty$-transitive or (P)-closed. Then, for every $t\in T$ such that $(G,G_{t},G_{t})$ has polynomial double-coset growth, we have
@@
The behaviour at $s=-1$ and the weighted Ihara zeta function
The zeta function $\mathcal{Z}_{\Gamma,u\to u}(s)$ is not the only growth series that has been considered for a weighted graph $(\Gamma,\omega)$. Another relevant example is the weighted Ihara zeta function $Z_{(\Gamma,W)}(x)$ introduced by A. Deitmar [9] for every graph $\Gamma$ with transition weight $W$ (cf. Section 7.3 for a brief recap). In Section 7.3, we provide a canonical way to construct a transition weight $W_{(\Gamma,\omega)}$ on a finite graph $\Gamma$ starting from an edge weight $\omega$ (cf. Example 7.11). The main result of the section focuses on a finite connected non-empty graph $\Gamma$ and subgraphs $\Gamma_{1},\Gamma_{2}\subseteq\Gamma$ satisfying the following: $\Gamma=\Gamma_{1}\cup\Gamma_{2}$, $\Gamma_{1}\cap\Gamma_{2}$ is a $1$-segment graph with edge set $\{a,\bar{a}\}$, and $t(a)$ and $o(a)$ are terminal vertices of $\Gamma_{1}$ and $\Gamma_{2}$, respectively. The graphs $\Gamma_{1}$ and $\Gamma_{2}$ carry restricted edge weights $\omega_{1}$ and $\omega_{2}$ from $\omega$, respectively. Hence we prove the following.
@@
In the setting before, assume that $\omega(E\Gamma)\subseteq\mathbb{Z}_{\geq 2}$ and that $\omega(a)\geq 3$ or $\omega(\bar{a})\geq 3$, for every $a\in E\Gamma$. Then,
@@
Finally, in Section 7 we discuss the behaviour at $s=-1$ for the double-coset zeta functions studied in Sections 5.4 and 5.5. In particular, we provide connections with the Euler–Poincaré characteristic of the group (cf. Section 7.2) and the Ihara zeta function of a weighted graph (cf. Section 7.3).
Acknowledgements
The author would like to thank her PhD advisors Christopher Voll and Thomas Weigel for the continuous support and instructive inputs throughout the preparation of this paper. The author also would like to thank Ilaria Castellano, Jan Moritz Petschick, José Pedro Quintanilha and Karthika Rajeev for interesting conversations and several useful comments on an early draft of the paper, and Evgeny Khukhro and Roman Sauer for interesting questions that inspired the results in Section 6.
The author was supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 491392403 – TRR 358. The author is a member of Gruppo Nazionale per le Strutture Algebriche, Geometriche e le loro Applicazioni (GNSAGA), which is part of the Istituto Nazionale di Alta Matematica (INdAM).

### v2

# arXiv:2409.01860v2[math.GR] 02 Mar 2026

@@ Abstract.
We study the double-coset zeta functions for groups acting on trees, focusing on weakly locally $\infty$-transitive or (P)-closed actions. After giving a geometric characterisation of convergence for the defining series, we provide explicit determinant formulae for the relevant zeta functions in terms of local data of the action. Moreover, we prove that the evaluation of these zeta functions at 1-1 satisfies the expected identity with the Euler–Poincaré characteristic of the group. Our framework suggests the definition of a new zeta function counting weighted paths in graphs, which recovers some double-coset zeta functions considered in this paper as particular cases. The behaviour at 1-1 of the latter zeta function sheds light on a connection with the Ihara zeta function of a weighted graph introduced by A. Deitmar.
@@ 1. Introduction
Double cosets play a prominent role in multiple aspects of group theory and beyond. For instance, they are the building blocks of Hecke algebras. Regarded as collections of cosets, they describe spheres with respect to the Weyl distance in the building associated with a Bruhat decomposition of a group [1]. Strictly related objects, namely the suborbits, are also widely studied in permutation group theory. It is common to arrange the suborbit sizes – provided they are all finite – in a non-decreasing sequence $(a_{n})_{n\geq 1}$ and estimate, for example, the growth of $a_{n}$ as a function of $n$ (see, for instance, [4] or [28, §5]).
@@
If there is some $\alpha\in\mathbb{R}_{\geq 0}$ such that $a_{n}(G,H,K)=O(n^{\alpha})$ for every $n$, we say that $(G,H,K)$ has polynomial double-coset growth.
If $(G,H,K)$ has polynomial double-coset growth, then an indirect (but often more convenient) way to study the sequence $(a_{n}(G,H,K))_{n\geq 1}$ is through the Dirichlet series the sequence generates, that is,
@@
where $s$ is a complex variable. By [17, §I.1, Theorem 3], this series converges in some half-plane $\{s\in\mathbb{C}\mid\mathrm{Re}(s)>\alpha\}$ and the function it determines is called the double-coset zeta function of $(G,H,K)$.
Zeta functions initially arose in number theory, although nowadays they are also an established tool in studying groups, rings and algebras (see M. du Sautoy’s survey [26] for motivation). In group theory, the seminal work of F. Grunewald, D. Segal, G. Smith [16] initiated the study of numerous zeta functions associated to finitely generated nilpotent groups or profinite groups. Double-coset zeta functions have been one of the first instances of zeta functions in the class of totally disconnected locally compact (= t.d.l.c.) groups which are possibly neither discrete nor profinite. Their introduction was motivated by an interesting behaviour at s=1s=-1. I. Castellano, G. Chinello and T. Weigel [10] have provided examples of unimodular t.d.l.c. groups $G$ with a compact open subgroup $K\leq G$ for which the meromorphic continuation of $\zeta_{G,K,K}(s)$, evaluated at s=1s=-1, recovers the Euler–Poincaré characteristic $\widetilde{\chi}_{G}$ of $G$ (in the sense of [10, §5]). Namely, one has
@@
The connection between growth series and Euler–Poincaré characteristics is not an isolated phenomenon, and usually it gives rise to surprising connections between two apparently unrelated objects. The reader is referred, for instance, to [13, 15, 30] or the introduction of [10] for some remarkable examples. One of the main goals of the present paper is to prove that the Euler–Poincaré identity holds in two relevant classes of unimodular t.d.l.c. groups acting on trees. More generally, we present a systematic study of the double-coset zeta functions for groups (not necessarily t.d.l.c.) acting on trees, including convergence criteria and explicit formulae.
Convention 1.0.1.
In what follows, every graph $\Gamma=V\Gamma\sqcup E\Gamma$ is meant in the sense of J-P. Serre [27], and $V\Gamma$ and $E\Gamma$ denote the set of vertices and the set of edges of $\Gamma$, respectively (cf. Section 2.2).
Moreover, we assume that every group action on a tree is without edge inversions and the tree has at least one edge (cf. Section 3.1). Finally, the stabilisers of adjacent vertices are assumed to be incomparable with respect to the inclusion.
@@
[…]
The second property we consider is (P)-closedness, a property that stems from the more general concept of actions on trees with Tits’ property (P). The latter property has been introduced by J. Tits [32, §4.2] and is recalled in Section 3.3. A group action on a tree $(G,T)$ is (P)-closed if it has Tits’ property (P) and $G\leq\mathrm{Aut}(T)$ is closed with respect to the permutation topology. Tits’ property (P) play a central role in the theory of groups acting on trees and even beyond. For instance, one may use a (P)-closed action to produce and study properties of simple groups acting on trees (cf. for instance [32, Théorème 4.5], [2, Theorem 7.3][24, Theorem 1.8], and [7]).
We thank one of the anonymous referees to have brought the attention to the following clarification.
Remark 1.0.1.
Another (equivalent) way to define (P)-closed actions on trees is to require that the group $G\leq\mathrm{Aut}(T)$ of the relevant action $(G,T)$ equals its (P)-closure (or (P1)-closure). The latter is a suitably closed subgroup of $\mathrm{Aut}(T)$ that has been first introduced by C. Banks, M. Elder, and G. Willis [2, §3] for locally finite trees under the initial name of $1$-closure. The term 11-closure” is problematic (cf. the footnote in [24, §2.3]), so it has been successively replaced by “(P)-closure” (or “(P1)-closure”, as there is a more general notion of “(Pk)-closure” for every $k\geq 1$).
In [24, §2.3], C. Reid and S. Smith extend the definition of (P)-closure to group actions on arbitrary trees. Their notion agrees with the one in [2] if the tree is locally finite, and is equivalent to the definition that we use in the present paper by [24, Theorem 2.18].
Remarkably, C. Reid and S. Smith [24] provide a complete classification of (P)-closed actions on trees by their local action diagram, which is a local datum attached to each group action on a tree. In Section 3.4 we briefly recall it. With this local description, one may study several global properties of the group action through more accessible features of the associated local action diagram. In the present paper, we exploit this approach more than once.
The reader may find it convenient to keep the following example in mind, as it represents a group action on a tree with both the two properties mentioned before.
@@
Let $\Gamma$ be a connected graph with a function $\omega\colon E\Gamma\to\mathbb{Z}_{\geq 1}$ (called edge weight). Following [8, §3.5], the pair $(\Gamma,\omega)$ admits an essentially unique universal cover $(T,\pi)$, which consists of a tree $T$ and a graph epimorphism $\pi\colon T\to\Gamma$ with the following property: for all $a\in E\Gamma$ and $v\in VT$ with $\pi(v)=o(a)$, the number of edges $e\in ET$ with origin $v$ and $\pi(e)=a$ is exactly $\omega(a)$. The group of deck transformations of $(\Gamma,\omega)$ is
@@
where $\mathrm{Aut}(T)$ denotes the group of automorphisms of $T$. The group $\mathrm{Aut}_{\pi}(T)$ is t.d.l.c. with respect to the permutation topology. Moreover, the $\mathrm{Aut}_{\pi}(T)$-action on $T$ is both weakly locally $\infty$-transitive and (P)-closed (cf. Example 3.12(i) and [32, §4.2]).
@@
The first step towards the study of double-coset zeta functions is to determine under which conditions a triple $(G,H,K)$ has the double-coset property or polynomial double-coset growth. Here we only consider the case of $G$ acting on a locally finite tree and we choose $H,K$ among the stabilisers of vertices or edges of $T$ (written $H=G_{t_{1}}$ and $K=G_{t_{2}}$, for $t_{1},t_{2}\in T=VT\sqcup ET$). In this setting, we provide the following characterisation.
Theorem A (cf. Theorems 5.8 and 5.9).
Let $G$ be a group that acts on a locally finite tree $T$ as in Convention 1.0.1 and with finite quotient graph. Assume that the action $(G,T)$ is weakly locally $\infty$-transitive or (P)-closed. Then the following are equivalent, for all $t_{1},t_{2}\in T$:
@@
    for every $e\in ET$, $G_{e}$ is not contained in any end stabiliser.
By Theorem A, both the double-coset property and the polynomial double-coset growth are independent of the choice of $t_{1}$ and $t_{2}$. A similar independence has been shown in [10, Proposition 6.2].
In Theorem A, if $(G,T)$ is (P)-closed then (iii) is equivalent to the condition that no end stabilisers are open or, equivalently, that no ends of $T$ are locally invariant (cf. [24, Definition 4.12, Proposition 4.23]).
A partial version of Theorem A with more general hypotheses on the action holds and can be found in Proposition 5.7.
@@
In the weakly locally $\infty$-transitive case, counting the $(H,K)$-double cosets is rephrased in a more accessible counting of certain paths in the quotient graph $\Gamma$ with a suitably defined weight (cf. Section 3.2). This also suggests the definition of a more general Dirichlet series $\mathcal{Z}_{\Gamma,u_{1}\to u_{2}}(s)$ associated to an arbitrary graph $\Gamma$ with edge weight $\omega\colon E\Gamma\to\mathbb{Z}_{\geq 1}$ and to $u_{1},u_{2}\in\Gamma$ (cf. Definition 5.11). The series $\mathcal{Z}_{\Gamma,u_{1}\to u_{2}}(s)$ recovers $\zeta_{G,G_{t_{1}},G_{t_{2}}}(s)$ whenever $G$ is a group acting weakly locally $\infty$-transitively on a locally finite tree $T$ with quotient graph $\Gamma$ and standard edge weight $\omega$ (cf. Section 3.2), and $t_{1},t_{2}\in T$ satisfy $G\cdot t_{1}=u_{1}$ and $G\cdot t_{2}=u_{2}$.
@@
In both cases, we can adapt ideas and techniques from graph theory (e.g., counting paths in graphs by their weight, see for instance [11] or [12]) and provide explicit formulae for $\zeta_{G,G_{t_{1}},G_{t_{2}}}(s)$, for all $t_{1},t_{2}\in T$.
@@
Theorem B (cf. Theorems 5.14 and 5.21).
Let $(G,T)$ be a group action on a locally finite tree as in Convention 1.0.1 that is weakly locally $\infty$-transitive or (P)-closed. Let $t_{1},t_{2}\in T$ be such that $(G,G_{t_{1}},G_{t_{2}})$ has polynomial double-coset growth. Then
@@
for explicitly defined square matrices $\mathcal{E}^{\bullet}(s)$ and $\mathcal{U}^{\bullet}_{t_{1},t_{2}}(s)$ whose entries are entire functions in $s\in\mathbb{C}$, and for a determined integer $\epsilon^{\bullet}_{t_{1}}(t_{2})$. Here $I^{\bullet}$ denotes the identity matrix of the same dimension as $\mathcal{E}^{\bullet}(s)$. In particular, $\zeta^{\bullet}_{G,G_{t_{1}},G_{t_{2}}}(s)$ extends to a meromorphic function on $\mathbb{C}$.
In Theorem B, the matrix $\mathcal{E}^{\bullet}(s)$ can be interpreted as a weighted adjacency matrix of the local structure in which we count the paths (cf. Definitions 5.12 and 5.19). The matrix $\mathcal{U}_{t_{1},t_{2}}^{\bullet}(s)$ and the integer $\epsilon^{\bullet}_{t_{1}}(t_{2})$ can be regarded as “perturbation data” given by the choice of $t_{1}$ and $t_{2}$.
@@
The behaviour at s=1s=-1 and the Euler–Poincaré characteristic
One of the main goals of the paper is the study of the local behaviour at s=1s=-1 of the relevant double-coset zeta functions. Unless it is a finite sum, the Dirichlet series in (1.2) does not converge at s=1s=-1. This underlines the importance of having a continuation of $\zeta_{G,H,K}(s)$ at least to s=1s=-1 (which in our context is provided by Theorem B) to carry out such an evaluation.
Addressing [10, Question G(b)], we prove that the Euler–Poincaré identity is also attained in our framework. A crucial step towards this goal is to reduce the evaluation at s=1s=-1 to the more accessible weakly locally $\infty$-transitive case (cf. Lemma 7.10). More specifically, we can only focus on $\mathcal{Z}_{\Gamma,u\to u}(-1)^{-1}$ for a finite connected edge-weighted graph $(\Gamma,\omega)$. After introducing suitable notions of unimodularity and Euler–Poincaré characteristic $\chi(\Gamma,u)$ at $u\in\Gamma$ on $(\Gamma,\omega)$ (cf. Definition 7.4), we deduce the following.
@@
In Theorem E, the hypothesis on $\omega$ guarantees that a formula analogous to the one in Theorem B is applicable to $\mathcal{Z}_{\Gamma,u\to u}(s)$, for all $u\in\Gamma$. The proof of Theorem E is based on some splitting formulae satisfied by $\mathcal{Z}_{\Gamma,u\to u}(s)^{-1}$, where $u\in\Gamma$, that we discuss in Section 6. The assumption that $\Gamma$ has no cycles of length $\geq 2$ guarantees their general applicability. By Theorem E and Lemma 7.10, we deduce the following.
@@
Let $G$ be a unimodular t.d.l.c. group acting on a locally finite tree $T$ as in Convention 1.0.1 and with compact open vertex stabilisers. Assume that the quotient graph $\Gamma$ and the associated standard edge-weight $\omega\colon E\Gamma\to\mathbb{Z}_{\geq 2}$ (cf. Section 3.2) satisfy the hypotheses of Theorem E. Suppose also that $(G,T)$ is weakly locally $\infty$-transitive or (P)-closed. Then, for every $t\in T$ such that $(G,G_{t},G_{t})$ has polynomial double-coset growth, we have
@@
The behaviour at s=1s=-1 and the weighted Ihara zeta function
The zeta function $\mathcal{Z}_{\Gamma,u\to u}(s)$ is not the only growth series that has been considered in a graph-theoretical setting. Within Dirichlet series, we mention the zeta functions associated to graphs and hypergraphs introduced by T. Rossmann and C. Voll [25], used by them to study the enumeration of conjugacy classes of unipotent groups over finite fields. Within growth series, a well-known power series associated to a graph is its Ihara zeta function [19], which relates closed paths in the graph with the spectrum of its adjacency matrix. In the present note, we focus on a variation of the Ihara zeta function, namely the weighted Ihara zeta function. This is a power series $Z_{(\Gamma,W)}(x)$ introduced by A. Deitmar [11] for every graph $\Gamma$ with transition weight $W$ (cf. Section 7.3 for a brief recap). In Section 7.3, we provide a canonical way to construct a transition weight $W_{(\Gamma,\omega)}$ on a finite graph $\Gamma$ starting from an edge weight $\omega$ (cf. Example 7.11).
The main result of the section is the following.
@@
Let $\Gamma$ be a finite connected non-empty graph, and $\Gamma_{1},\Gamma_{2}$ be subgraphs satisfying $\Gamma=\Gamma_{1}\cup\Gamma_{2}$, such that $\Gamma_{1}\cap\Gamma_{2}$ is a $1$-segment graph, i.e., a Serre graph with two distinct vertices and an edge couple $\{a,\bar{a}\}$ connecting them, such that $t(a)$ and $o(a)$ are leaves of $\Gamma_{1}$ and $\Gamma_{2}$, respectively. Let also $\omega\colon E\Gamma\to\mathbb{Z}_{\geq 2}$ satisfy $\omega(a)\geq 3$ or $\omega(\bar{a})\geq 3$ for every $a\in E\Gamma$. Then,
@@
Relating Dirichlet series and power series, even if generated by a common sequence, is usually a non-trivial problem. In our case, the relation established in Theorem G has been facilitated by explicit determinant formulae for each of the two types of zeta functions involved. Theorem G sheds light on the behaviour of the weighted Ihara zeta function $Z_{(\Gamma,W)}(t)$ at $t=1$ which, to the best of the author’s knowledge, has been never studied so far. In the same spirit as in Theorem E, it would be interesting to find an explicit expression of $Z_{(\Gamma,W)}(1)$ in terms of suitable invariants of $(\Gamma,W)$.
@@
Finally, in Section 7 we discuss the behaviour at s=1s=-1 for the double-coset zeta functions studied in Sections 5.4 and $5.$ In particular, we provide connections with the Euler–Poincaré characteristic of the group (cf. Section 7.2) and the Ihara zeta function of a weighted graph (cf. Section 7.3).
Acknowledgments
The author thanks her Ph.D. advisors Christopher Voll and Thomas Weigel for continuous support and instructive input throughout the preparation of this paper. She also thanks Ilaria Castellano, Jan Moritz Petschick, José Pedro Quintanilha, and Karthika Rajeev for interesting conversations and several useful comments on an early draught of the article, and Evgenii Khukhro and Roman Sauer for interesting questions that inspired the results in Section 6. Finally, she is grateful to the two anonymous referees for valuable and extremely instructive suggestions that improved the article.
The author is a postdoctoral researcher of the Fonds de la Recherche Scientifique – FNRS. The author was supported by DFG TRR 358 –- Project-ID 491392403, and RTG 2229 “Asymptotic Invariants and Limits of Groups and Spaces”. She is also a member of GNSAGA, as a part of Istituto Nazionale di Alta Matematica (INdAM).